At current exchange rates it takes 0.1475 Canadian dollars to buy Chinese yuan (CNY), and 0.6052 Canadian dollars to buy one Brazilian real (BRL). What must the yuan/real exchange rate be in order to eliminate arbitrage opportunities?
Answer #1

1 CNY = 0.1475 Canadian dollars, it is expressed as : CAD/CNY = 0.1475

1 BRL = 0.6052 Canadian dollars, it is expressed as : CAD/BRL = 0.6052

yuan/real exchange rate = CNY/BRL = (CNY/CAD)/(BRL/CAD) = 1/(CAD/CNY) X (CAD/BRL)

yuan/real exchange rate = 1/0.1475 X 0.6052 = 4.1031 CNY/ BRL

ANSWER : 4.1031 CNY/ BRL (Thumbs up please)